suppose the following data describe a nation’s population: Year 1 Year 2 Population 200 million 203 million Labor Force 120 million 125 million Unemployment rate 6 percent 6 percent a. How many people are unemployed in each year? b. How many people are employed in each year? c. Compute the employment rate (i.e., number employed/population) each year. d. How can the employment rate rise when the unemployment rate is constant?


1
Expert's answer
2021-04-26T19:30:14-0400

a,b) unemployed people in Year 1: "120 \times 0,06 = 4,8 million"

employed people in Year 1: "120 - 4,8 = 117,2million"

unemployed people in Year 2: "125 \times 0,06 = 7,25 million"

employed people in Year 2: "125 - 7,25 = 118,75 million"

c) employment rate Year 1 and 2: "100\% - 6\% = 94\%"

d) it is possible when the labor force and number of employed people rises while the pace of employed people rise must be higher